• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 09:58:53 on January24,2002

    Entry number 80554

    keyword=target situation

    I was paged in this morning for two probelms:
    1) Alarm handler not working.
    2) Lifter indicator was not working while in LH2 no rad position
    during the night shift.
    It was found that ioc was dead. So rebooted ioc. After rebooting, the target
    was found in LH2 no rad position! (While before reboot, it was in optics target
    position). I moved the target to optics again, it seems fine. Re-entered all the
    alarm set values. The save alerm to default file not working yet. Informed
    Chris Keith about the alarm default issue and the lift moving by it self due to
    rebooting. It should not happen and possiblly somebody clicked the LH2 no-rad
    position while ioc died (though checking with the target operator, nobody clicked).
    We need to be careful with the lifter and try to minimize motion.
    Also asked Chris to change target vertical encoder by 2 mm (target
    was high by 3.5 mm). When he finishes, we will need to reboot to have the new
    values in effect. Also the high power heater default set point is 22 K for loop 3,
    Chris will work to let it take the modified value (19K).
    For now, after reboot, one needs to check the HP heater PID set value and change
    to 19K if it is not 19K already. Also need to manually enter all teh appropriate alarm
    limits values.

    IOC dead one more time by itself, and rebooted.